# Insights into DeepSeek-V3: Scaling Challenges and Reflections on Hardware for AI Architectures

### v1

@@ 6.4. Discussion on Memory-Semantic Communication and Ordering Issue
Several approaches are possible. For instance, the receiver could buffer atomic messages and use packet sequence numbers for in-order processing. However, an acquire/release mechanism is both more elegant and efficient. We suggest a simple conceptual mechanism, Region Acquire/Release (RAR) mechanism, wherein receiver hardware maintains a bitmap to track the state of the RNR memory region, and acquire/release operations are scoped to the RAR address range. With minimal bitmap overhead, this enables efficient, hardware-enforced ordering, eliminating explicit sender-side fences and delegating ordering to hardware—ideally on the NIC or I/O die. Importantly, the RAR mechanism benefits not only memory-semantic operations but also message-semantic RDMA primitives, thus broadening its practical applicability.
@@ 6.6.2. Suggestions:
  • DRAM-Stacked Accelerators: Leveraging advanced 3D stacking technologies, DRAM dies can be vertically integrated atop a logic die, thereby enabling exceptionally high memory bandwidth, ultra-low latency, and a practical memory capacity (though stack-limited). This architectural paradigm proves remarkably advantageous for ultra-fast inference in MoE models, where memory throughput is a critical bottleneck. Architectures such as SeDRAM(Wang et al., 2023) exemplify the potential of this approach, delivering unprecedented performance for memory-bound workloads.
  • System-on-Wafer (SoW): Wafer-scale integration (Lie, 2022) can maximize computational density and memory bandwidth, addressing the needs of ultra-large-scale models.

### v2

@@ 6.4. Discussion on Memory-Semantic Communication and Ordering Issue
Several approaches are possible. For example, the receiver can buffer atomic messages and use packet sequence numbers (PSN) to ensure in-order processing; this method is straightforward and effective for maintaining correctness. Alternatively, a region-based acquire/release (RAR) mechanism is particularly attractive, as it enables true acquire/release semantics for remote memory access and offers greater flexibility for a range of workloads. In this approach, the receiver hardware maintains lightweight metadata—such as bitmaps or region-based counters—to track the state of memory regions. Acquire and release operations are scoped to specific address ranges, enabling efficient, hardware-enforced ordering without explicit sender-side fences. Notably, both approaches are amenable to implementation on the NIC or I/O die and are applicable to both memory-semantic and message-semantic RDMA primitives, thereby broadening their practical utility.
@@ 6.6.2. Suggestions:
  • DRAM-Stacked Accelerators: Leveraging advanced 3D stacking technologies, DRAM dies can be vertically integrated atop a logic die, thereby enabling exceptionally high memory bandwidth, ultra-low latency, and a practical memory capacity (though stack-limited). This architectural paradigm proves remarkably advantageous for ultra-fast inference in MoE models, where memory throughput is a critical bottleneck. Architectures such as SeDRAM(10185427) exemplify the potential of this approach, delivering unprecedented performance for memory-bound workloads.
  • System-on-Wafer (SoW): Wafer-scale integration (cerebras) can maximize computational density and memory bandwidth, addressing the needs of ultra-large-scale models.